OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14 2 1 -1 -2

OpenStudy (loser66):

\[\huge 4x+3y=10\\ -\\\huge4x+5y=14\\-----------------\\=?\]

OpenStudy (loser66):

yes, I want you take the first equation minus the second one, what do you get?

OpenStudy (anonymous):

0 -2y = -4

OpenStudy (loser66):

so y=?

OpenStudy (anonymous):

0.5?

OpenStudy (loser66):

nope, divided both sides by -2

OpenStudy (anonymous):

2?

OpenStudy (loser66):

bingo, you have y =2 , replace to either of the equations, I pick the first one, now, replace !! what do you have?

OpenStudy (anonymous):

x = 1

OpenStudy (loser66):

you are right, I don't know on which way you get that, but YOU ARE RIIIIIGHT

OpenStudy (loser66):

know how to get, right? hopefully you say YES

OpenStudy (anonymous):

4x + 3(2) = 10 4x + 6 = 10 4x + 6 - 6 = 10 - 6 4x = 4 4x/4 = 4/4 x = 1

OpenStudy (loser66):

BINGO!!!

OpenStudy (anonymous):

Cheers! Thanks a million mate!

OpenStudy (loser66):

yw

OpenStudy (anonymous):

=)